A boat on the Hudson River travels the same distance in 1/2 hour downstream as it can going upstream for an hour. If the boat travels 12 mph in still water, what is the speed of the current?
Please help! There are no examples like this one in the book.
Thanks!

Answer by stanbon(75887)

A boat on the Hudson River travels the same distance in 1/2 hour downstream as it can going upstream for an hour. If the boat travels 12 mph in still water, what is the speed of the current?
-------------------------------------
Downstream DATA:
time = 1/2 hr ; rate = (12+c) mph ; distance = rt = (1/2)(12+c) = 6 + c/2 miles
--------------------------------------
Upstream DATA:
time = 1 hr ; rate = (12-c) mph ; distance = rt = 1(12-c) mph
--------------------------------------
Equation:
distance up = distance down
6 + c/2 = 12 - c
(3/2)c = 6
c = 4 mph (speed of the current)
====================================
Cheers,
Stan H.